OpenStudy (anonymous):

in my homework I am asked to solve for h in this equation d=1.2square root symbol with an h (sorry I do not know how to show symbol)

OpenStudy (anonymous):

\[ d /\sqrt{h}=1.2\] are you looking for smntng like this? im not sure what your values are for d and h

OpenStudy (anonymous):

oh i meant d/1.2=sqr of h then just square the other side, so d^2/1.2^2=h